OpenStudy (bananas1234):

Which equation is the equation of a 10cm radius circle with it's center at (3, -6)

Mehek (mehek14):

\((x-h)^2+(y-k)^2=r^2\) H is the x value of the coordinate, K the y value, and R = radius

OpenStudy (bananas1234):

so(X - 3)^2 + (y - -6)^2 = r^2

OpenStudy (bananas1234):

but the two negatives would turn into a positive

Mehek (mehek14):

Close (x-3)^2 is correct (y - - 6)^2 should become (y + 6)^2

Mehek (mehek14):

And for the r^2, you would but 10^2 = ?

OpenStudy (bananas1234):

20

Mehek (mehek14):

no 10^2 = 10 * 10 not 10 * 2

OpenStudy (bananas1234):

oh sorry 100

Mehek (mehek14):

Ok so the final answer is: \((x-3)^2+(y+6)^2=100\)

OpenStudy (bananas1234):

Tank you for your help!!

OpenStudy (bananas1234):

thank*

Mehek (mehek14):

You're welcome :)